# Inference Scaling fLaws: The Limits of LLM Resampling with Imperfect Verifiers

### v2

@@ A.2.2 Additional details on excluded tasks from MBPP+
A.2.3 Additional details on excluded tasks from HumanEval+
@@
A.3 Additional details on Section 4
In addition to the empirical analysis presented in Section 4, in this appendix, we provide a theoretical model that formalizes the limitations of inference scaling with imperfect verifiers and generalizes our findings to other benchmarks. We build on the verifier-based judge setup introduced by Davis et al. [16]. We provide a Python notebook with the implementation of our model [^3] See: https://colab.research.google.com/drive/13k81mNLcPSH5VAjUtwBGMncS8lfAaTx3?usp=sharing.
Figure 14: Even with zero computational cost, the optimal number of samples is finite and very low ($K\leq 3$). For this plot, we set the parameters as empirically observed for Llama 3.1 8B on HumanEval (see Table 3 for the exact values). The left plot shows the expected value of generating additional candidate solutions as a function of the number of attempts $K$ for various cost-benefit ratios. For all cost-benefit ratios, the expected value peaks at very low $K$, after which it begins to decline, indicating negative returns from additional sampling. The right depicts the probabilities of generating a correct solution vs. a false positive at each step $K$. As $K$ increases, the likelihood of generating a correct solution decreases, while the probability of generating a false positive increases. There is a trade-off between continued sampling and increasing risk, emphasizing the limitations of scaling inference compute with imperfect verifiers. Note that when setting the cost of a false positive to be 10 times higher than the benefit of a true positive, the optimal number of samples becomes $K=0$ (Fig. 15).
@@ A.3 Additional details on Section 4
Parameter values used in model underlying Fig. 14
@@
Table 3: Parameter values used in the model setup in Fig. 14 as observed for the Llama 3.1 8B model evaluated on HumanEval+. These values reflect the empirically observed probabilities and prior settings. Following the observed empirical task difficulty distribution as shown in Fig. 6, in this setup we assume tasks with Pass@1 $\geq 0.5$ to be easy, and those with Pass@1 $<0.5$ to be hard.
@@
After each rejection, the belief that the task is of type $T_{2}$ (hard) increases. The posterior probability that the task is of type $T_{1}$ or $T_{2}$ after $k-1$ rejections is:
@@
As more rejections occur, it usually becomes more likely that the task is hard ($T_{2}$). In Fig. 14, we see how the belief that the task is easy decreases, while the belief that the task is hard increases as the number of attempts $K$ grows.
@@
The belief-weighted probability of returning a correct or false positive at attempt $k$, conditional on the $k-1$ previous attempts being rejected are:
@@
In Fig. 14, the right plot shows the evolution of $P_{\text{TP}}^{(k)}$ and $P_{\text{FP}}^{(k)}$ as the number of attempts $K$ increases. Initially, the probability of generating a correct solution is higher, but for higher $K$, the probability of generating a false positive increases.
@@
The optimal number of attempts, $K_{\text{opt}}$, is the value of $K$ that maximizes the reward, which are shown across models and for various $V_{FP}/V_{TP}$-ratios in Fig. 15.
A.4 Inference scaling curve for GPT-4o
Figure 16: Inference scaling curves in the presence of a cost for GPT-4o. In addition to the models in Fig. 1, we provide the inference scaling curves for GPT-4o as the model with the highest single-sample accuracy on on HumanEval+ in our experiments. We find that the benefits of search are minimal (i.e. curves are flat) in line with what we expect from the empirical task difficulty distribution shown in Fig. 10.
A.5 Additional details on Section 5
@@
A.5.1 Details on data and implementation
We used the implementation provided by Zheng et al. [79] to collect samples and evaluate the different code readability metrics. Each code quality metric had a separate prompt instructing the model to follow certain guidelines (Section A.5.2). For each model and code quality instruction, we generated 50 samples per task on HumanEval+. As for our main experiments, we set the temperature to 0.8. All other parameters were set to their default value as provided with the implementation. [^4] See: https://github.com/jszheng21/RACE/tree/3b8ee591abd5febd8ae8ec17c7b9907949c5e1d5
A.5.2 Prompt examples for readability metrics
1) Naming conventions
@@ A.5.2 Prompt examples for readability metrics
[…]
2) Code length
[…]
3) Commenting guidelines
[…]
Figure 18: Prompt templates for each metric of code readability we consider in our experiments for Section 5 following Zheng et al. [79].
A.5.3 Qualitative examples of false positives
@@
(a) Example implementation generated by GPT-3.5 which fails additional unit tests from MBPP+.
@@
(b) Standard unit tests for the MBPP/7 task.
@@
4 new_lst = [word for word in lst if len(word) %
@@
Figure 27: MBPP/565 asks to split a string into characters. The implementation generated by Command-Light incorrectly filters for alphabetic characters only and fails to handle empty strings correctly.
@@
Appendix B Reproducibility statement
We release code to reproduce all experimental results of this paper in a GitHub repository [^5] See: https://github.com/benediktstroebl/inference-scaling-limits. This repository also contains all code samples for all models used in our experiments. We also provide an implementation of the theoretical model in Section A.3 as a Python notebook [^6] See: https://colab.research.google.com/drive/13k81mNLcPSH5VAjUtwBGMncS8lfAaTx3?usp=sharing.

### v3

@@ B.2 Additional details on excluded tasks from MBPP+
B.3 Additional details on excluded tasks from HumanEval+
@@
Appendix C Additional details on Section˜4
In addition to the empirical analysis presented in Section˜4, in this appendix, we provide a theoretical model that formalizes the limitations of inference scaling with imperfect verifiers and generalizes our findings to other benchmarks. We build on the verifier-based judge setup introduced by Davis et al. (2024). We provide a Python notebook with the implementation of our model [^4] See supplementary materials, file limits_to_inference_scaling_model.ipynb.
Figure 13: Even with zero computational cost, the optimal number of samples is finite and very low ($K\leq 3$). For this plot, we set the parameters as empirically observed for Llama 3.1 8B on HumanEval (see Table˜3 for the exact values). The left plot shows the expected value of generating additional candidate solutions as a function of the number of attempts $K$ for various cost-benefit ratios. For all cost-benefit ratios, the expected value peaks at very low $K$, after which it begins to decline, indicating negative returns from additional sampling. The right depicts the probabilities of generating a correct solution vs. a false positive at each step $K$. As $K$ increases, the likelihood of generating a correct solution decreases, while the probability of generating a false positive increases. There is a trade-off between continued sampling and increasing risk, emphasizing the limitations of scaling inference compute with imperfect verifiers. Note that when setting the cost of a false positive to be 10 times higher than the benefit of a true positive, the optimal number of samples becomes $K=0$ (Figure˜14).
@@ C.1 Model setup
Parameter values used in model underlying Figure˜13
@@
Table 3: Parameter values used in the model setup in Figure˜13 as observed for the Llama 3.1 8B model evaluated on HumanEval+. These values reflect the empirically observed probabilities and prior settings. Following the observed empirical task difficulty distribution as shown in Figure˜6, in this setup we assume tasks with Pass@1 $\geq 0.5$ to be easy, and those with Pass@1 <0.5<0.5 to be hard.
@@
After each rejection, the belief that the task is of type $T_{2}$ (hard) increases. The posterior probability that the task is of type $T_{1}$ or $T_{2}$ after k1k-1 rejections is:
@@
As more rejections occur, it usually becomes more likely that the task is hard ($T_{2}$). In Figure˜13, we see how the belief that the task is easy decreases, while the belief that the task is hard increases as the number of attempts $K$ grows.
@@
The belief-weighted probability of returning a correct or false positive at attempt $k$, conditional on the k1k-1 previous attempts being rejected are:
@@
In Figure˜13, the right plot shows the evolution of $P_{\text{TP}}^{(k)}$ and $P_{\text{FP}}^{(k)}$ as the number of attempts $K$ increases. Initially, the probability of generating a correct solution is higher, but for higher $K$, the probability of generating a false positive increases.
@@
The optimal number of attempts, $K_{\text{opt}}$, is the value of $K$ that maximizes the reward, which are shown across models and for various $V_{FP}/V_{TP}$-ratios in Figure˜14.
C.2 Inference scaling curve for GPT-4o
Figure 15: Inference scaling curves in the presence of a cost for GPT-4o. In addition to the models in LABEL:fig:empirical_scaling_curves, we provide the inference scaling curves for GPT-4o as the model with the highest single-sample accuracy on on HumanEval+ in our experiments. We find that the benefits of search are minimal (i.e. curves are flat) in line with what we expect from the empirical task difficulty distribution shown in Figure˜9.
Appendix D Additional details on Section˜5
@@
D.1 Details on data and implementation
We used the implementation provided by Zheng et al. (2024) to collect samples and evaluate the different code readability metrics. Each code quality metric had a separate prompt instructing the model to follow certain guidelines (Section˜D.2). For each model and code quality instruction, we generated 50 samples per task on HumanEval+. As for our main experiments, we set the temperature to 0.8. All other parameters were set to their default value as provided with the implementation. [^5] See: https://github.com/jszheng21/RACE/tree/3b8ee591abd5febd8ae8ec17c7b9907949c5e1d5
D.2 Prompt examples for readability metrics
@@
Figure 17: Prompt templates for each metric of code readability we consider in our experiments for Section˜5 following Zheng et al. (2024).
D.3 Qualitative examples of false positives
1def get_positive(l: list):
2 return [int(item) for item in l if item > 0]
a) Example implementation generated by CodeGen-6B which fails additional unit tests from HumanEval+.
1# Standard unit tests
3assert get_positive([-1, 2, -4, 5, 6]) == [2, 5, 6]
4
5# Example unit test from extended test suite
b) Standard unit tests for the HumanEval/30 task and one example test from the extended test suite of HumanEval+.
Figure 18: Example implementation that passes the standard tests but is incorrect. In HumanEval/30, the model is asked to Return only positive numbers in the list. The code was generated by CodeGen-6B (Nijkamp et al., 2023) and passes the tests provided with the original HumanEval benchmark but fails the extended tests. The implementation contains a logical error due to unnecessarily converting positive floats to integers.
@@ D.3 Qualitative examples of false positives
Example implementation generated by GPT-3.5 which fails additional unit tests from MBPP+.
@@
Standard unit tests for the MBPP/7 task.
@@
4 new_lst = [word for word in lst if len(word) % 2 == 0]
@@
Figure 27: MBPP/565 asks to split a string into characters. The implementation generated by Command Light incorrectly filters for alphabetic characters only and fails to handle empty strings correctly.
@@
Appendix E Declaration of LLM usage and compute resources
In this work, LLMs are the main subject of study and used for sampling solutions for coding benchmarks. For all our experiments using OpenAI models, we utilized the endpoints provided by OpenAI, either directly or through the Azure OpenAI Service. For the analysis on HotPotQA using Llama models, we relied on the endpoints provided by Together.ai. As our work primarily relied on external APIs, we did not use any GPUs for inference and our experiments did not require training of LLMs.
Appendix F Impact Statement
Our work aims to enhance the technical understanding of the limitations of inference scaling methods. In particular, we contribute new findings on resampling in combination with imperfect verifiers and how this approach can fail to lift the accuracy of weaker models to match the performance of stronger models. While these findings have implications for the training of reasoning models and the deployment of compound AI systems—including code generation tools—their broader ethical and societal implications mirror those already familiar in the development and use of large-scale language models. We do not identify any additional, domain-specific concerns that arise uniquely from our study. Instead, our results reinforce the importance of reliable evaluation metrics and thorough verification methods, which in turn support safer and more trustworthy applications in coding as well as in other areas where compound AI systems are increasingly adopted.